A plane of mass 1500 kg takes off with a speed of 200 km/h at the end of a runway 1 km long. If it starts from rest at the other end of the runway, find the average force being exerted by the propellers during takeoff (neglect friction).


1
Expert's answer
2021-07-08T13:28:20-0400

"F=ma"


"s=\frac{v^2}{2a}\to a=\frac{v^2}{2s}=\frac{55.56^2}{2\cdot1000}=1.54\ (m/s^2)"


"F=ma=1500\cdot1.54=2310\ (N)" . Answer